Assuming the following information calculate the acid (quick) ratio: current liabilities $25,000. long term liabilities $50,000, cash $10,000, inventory $30,000, accounts receivable $30,000.
Answer #1

Quick assets=cash+AR

=10,000+30,000

=$40,000

Quick ratio=Quick assets/Current liabilities

=40,000/25000

=1.6